OpenStudy (anonymous):

You are selling tickets for a high school basketball game. Student tickets cost $3 and general admission tickets cost $5. You sell 350 tickets and collect $1,450. How many student tickets did you sell?

jigglypuff314 (jigglypuff314):

3x + 5y = 1450 x + y = 350

OpenStudy (anonymous):

But how do I solve it?

jigglypuff314 (jigglypuff314):

I used substitution method :) x + y = 350 becomes y = -x + 350 which you can plug into 3x + 5y = 1450 so 3x + 5(-x+350) = 1450 distribute and solve for x

OpenStudy (anonymous):

First, i distribute. Which would be 3x-5x+350=1450?

jigglypuff314 (jigglypuff314):

close :) don't forget that the 5 gets multiplied to the 350 too so 3x - 5x + 1750 = 1450 then add like terms and solve for x :)

OpenStudy (anonymous):

2x=-300

OpenStudy (anonymous):

X=150

jigglypuff314 (jigglypuff314):

Correct! :)

OpenStudy (anonymous):

negative or positive 150?

jigglypuff314 (jigglypuff314):

I don't think you can have a negative number of tickets :P if you were wondering why... 3x - 5x = -2x so -2x = -300 divide both sides by -2 negative divided by negative= positive so x = positive 150

OpenStudy (anonymous):

True that and that makes sense. Thank you.

jigglypuff314 (jigglypuff314):

you're welcome :)